OpenStudy (anonymous):

the perimeter of a rectangle is twice the sum of its length and width. The perimeter is 34 meters and its length is 2 meters more than twice its width. What is its length?

OpenStudy (anonymous):

length = 5

OpenStudy (anonymous):

thankyou Amalu... can you possibly explain how you get that answer so next time I'll know. I appreciate your knowledge.

OpenStudy (anonymous):

perimeter = 2(l+b) 34=2(x+2+2x) 34=2x+4+4x 34=6x+4 6x=30 x=30/6 x=5

OpenStudy (anonymous):

34 = 2 ((w + 2) + w), w=15/2=7.5 l=w+2=19/2=9.5 2(9.5+7.5)=34

OpenStudy (anonymous):

remember... i have the mentality of a 5yr old. (ha ha) and need a basic explanation... thank you again